[Brahms-l] Shift report 20050529 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 29 2005 - 23:59:43 EDT
Shifter leader was ejk
Shifters were ps


1. two stores. 
   First  : dumped by high background at STAR
   
2. 15063 - 15069 : data
   15072 - 15110 : daq test, debugging run
   15111 : data

3. We had DAQ problem, and JH fixed.
   
    

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 29 23:59:56 2005

This archive was generated by hypermail 2.1.8 : Mon May 30 2005 - 00:00:09 EDT